<div dir="ltr"><div class="gmail_default" style="font-family:verdana,sans-serif">Dear all,</div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Apologies for losing the audio connection in the meeting today, it was quite heated and I didn't see that the phone switched off from speak-mode -- I will summarise the main points below.</div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">During controlled access today, while the accelerator vacuum leak was being repaired:</div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">* Krister worked on the over-heating connector lead on the power supply for the Moller magnet. He has disconnected it, cleaned it, applied a special sort of lubricant (details in his logbook entry, will appear tomorrow morning) and tested it by switching the magnet on for a few hours. The temperature crept up slightly, but both leads seemed to be behaving equally, which is a good sign. Krister will test it more extensively tomorrow morning.</div><div class="gmail_default" style="font-family:verdana,sans-serif">* Rafo adjusted FC offset (from ~40 to 250 Hz).</div><div class="gmail_default" style="font-family:verdana,sans-serif">* Florian tried to debug discriminator channel in slot 20 channel 2 of BAND, which was giving a slightly lower rate: unplugging and replugging didn't change that, so it'll be kept as it is. </div><div class="gmail_default" style="font-family:verdana,sans-serif">* Raffaella replaced a broken cable on the missing FT Hodoscope channel and saw a signal on fADC -- issue appears fixed: <a href="https://logbooks.jlab.org/entry/3750244" style="font-family:Arial,Helvetica,sans-serif" target="_blank">https://logbooks.jlab.org/entry/3750244</a></div><div class="gmail_default" style="font-family:verdana,sans-serif">* Yuri is looking into possible relation between radiation levels and the leakage current on SVT. For the moment inconclusive, but he would like to install an extra gamma monitor between SVT and the solenoid.</div><div class="gmail_default" style="font-family:verdana,sans-serif">* Raffaella and I unplugged and replugged HV cable and fADC connector for the CND channel S14L1R, which has been occasionally tripping off. It seems responsive, but needs to be watched.</div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Raffaella has checked the RF offset for the recent runs and it looks stable: <a href="https://logbooks.jlab.org/entry/3750178" style="font-family:Arial,Helvetica,sans-serif">https://logbooks.jlab.org/entry/3750178</a></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Nick Z has completed gain-matching of the FT Hodoscope and uploaded the new constants, which will be implemented from the next run: <a href="https://logbooks.jlab.org/entry/3750212" style="font-family:Arial,Helvetica,sans-serif" target="_blank">https://logbooks.jlab.org/entry/3750212</a></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Yordanka is producing monitoring histograms for the runs as they get cooked, they are now saved here: <a href="https://clasweb.jlab.org/clas12mon/rg-b/pass0/v17_fall2019/offline/">https://clasweb.jlab.org/clas12mon/rg-b/pass0/v17_fall2019/offline/</a></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Nick M. presented a study of HTCC response (please find his slides attached), where he compared the distributions between a spring run and one from now. There is a spike at low values in the Sping data, no spike in the current runs -- it was suggested that the difference is due to the new trigger parameters (addition of DC roads). He also pointed out that the distribution for S2 HS2 R1 is skewed to lower values and suggested that this can be corrected by calibrating to the peak position, rather than the mean of the distribution. Accuracy of the the gain matching for the HTCC channels is on the order of 5-10%, so the slights skew in that one channel was not considered to be a big issue.</div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Yuri S. gave a verbal update on the status of HTCC. The mirror manufacturers are still not providing info on the exact composition of the mirror coating (it's therefore hard to evaluate their aging). Near-future plans include replacement of quartz-window PMTs and of active dividers with passive dividers. New power supplies have also been ordered but not yet received.</div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Valery showed a 2D plot of the number of photoelectrons in each sector of the HTCC, from the current production data. Many of the half-sectors, especially in the outer ring, showed a structure where there were many more nphe in one half of the sector than in another. The half-sectors have a single mirror, so the sharp separation of these high- and low-nphe regions was decidedly odd. There was much heated discussion and argument over what might be causing this and whether this has been observed before (we lost audio connection in the scuffle). Valery called for HTCC simulations.</div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Valery then presented a plot of theta vs momentum for the electrons, which showed a large number of events at low theta. low momentum -- with a Q^2 below 0.5. He suggested that since they don't pose physics interest we could significantly reduce our electron trigger rate by cutting them out at the level of triggering (applying a Q^2 cut on DC roads). We agreed that he would produce new trigger firmware to do this and we would take a short (~half hour) run to test this. In the meantime we would continue taking data with the current trigger and only implement this once it's validated and if RG-B agrees. Current data rate (700 MB/s) is only sustainable if we're not taking data for about half the time, to meet the limits set for us. Should we get good, steady beam, our data rate will become a problem. </div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">About that -- the vacuum loss in the accelerator this morning was traced to a flange which was not properly closed (unclear why that was so or why it's only become a problem now -- mice?), they have tightened it and are pumping the vacuum down. This will take all night and it's not known whether it will fix the problem. Earliest beam delivery is expected tomorrow at noon. Swing and owl shifts are cancelled. </div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">The hall remains in Controlled Access and we have ramped down the solenoid and torus to 100A. If anyone needs access to the hall tonight, please contact me first -- the counting room will be unmanned. Access tomorrow morning will be as normal. </div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Planned work for the hall tomorrow: Krister will check Moller magnet again, Florian will get neutron counters for BAND removed, Yuri plans to install the extra gamma monitor he needs for SVT studies.</div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">A spin dance was planned for tomorrow (Halls A + C) -- when this eventually happens the consensus was that we do not participate but use the time to make a zero-field alignment measurement and take empty target or low luminosity data. If they change the Wien angle, we will then do a Moller measurement.</div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Have a good evening!</div><div class="gmail_default" style="font-family:verdana,sans-serif">Daria</div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">   </div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div></div><br><div class="gmail_quote"><div dir="ltr" class="gmail_attr">On Mon, 9 Dec 2019 at 11:40, Daria Sokhan <<a href="mailto:daria@jlab.org" target="_blank">daria@jlab.org</a>> wrote:<br></div><blockquote class="gmail_quote" style="margin:0px 0px 0px 0.8ex;border-left:1px solid rgb(204,204,204);padding-left:1ex"><div dir="ltr"><div dir="ltr"><div class="gmail_default" style="font-family:verdana,sans-serif">Dear all, </div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">We'll have the usual RC meeting upstairs in the counting house at 2pm. </div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Bluejeans:   <a href="https://urldefense.proofpoint.com/v2/url?u=https-3A__bluejeans.com_230805731&d=DwMFaQ&c=CJqEzB1piLOyyvZjb8YUQw&r=p0fOsYOqPes7kSrF0-ldLy8UQdwZWwpYV125WHZsrFE&m=bxMdN_MgodYZhFpK2RiAA3YZjwbYDgwEepbV0N-VDY4&s=0BPGwMe_soWKRhzk2lK-crzoQtg7vmiNMUWPakyPG00&e=" target="_blank">https://bluejeans.com/230805731</a></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">There has been another day of quiet physics running until ~5am this morning, when everyone returned to work and the accelerator threw a tantrum. MCC are currently down over vacuum problems.</div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Hall B is in controlled access at the moment -- anyone who needs to do anything down there, this is your chance! Current best MCC estimate for beam return is 1pm -- pre-scale by the usual factor. </div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif">Speak to you soon!</div><div class="gmail_default" style="font-family:verdana,sans-serif">Daria</div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif"><br></div><div class="gmail_default" style="font-family:verdana,sans-serif"> </div></div><div class="gmail_quote"><blockquote class="gmail_quote" style="margin:0px 0px 0px 0.8ex;border-left:1px solid rgb(204,204,204);padding-left:1ex">
</blockquote></div></div>
</blockquote></div>